ROTOR-VIBRATOR COUPLING EFFECTS IN METHYL CH STRETCHING OVERTONE SPECTRA OF GASEOUS TOLUENE

\underline{D. CAVAGNAT}, L. LESPADE and C. LAPOUGE, LPCM, Université Bordeaux I, F-33405 Talence Cedex.

The methyl CH stretching overtone spectra of toluene and of its deuterated derivatives (C_{6}D_{5}CH_{3}, C_{6}D_{5}CH_{2}D and C_{6}D_{5}CHD_{2}) have been recorded by FTIR (Δv=1 to 4) and by intracavity dye laser photoacoustic spectrometry (Δv=5 and 6) at low resolution (0.5 to 2 cm^{−1}). The different couplings likely to affect these spectra (anharmonic coupling of the νCH mode with the methyl group internal rotation, Fermi resonance coupling with combination states involving CH stretch/methyl bending modes) are considered. The parameters used to model these couplings are determined partly from ab-initio calculations^{a}, partly from fitting simultaneously the spectra of the different isotopic compounds. A preliminary analysis of the obtained results is presented. An attempt to compare them with those determined from the study of the corresponding overtone spectra of nitromethane ^{b}(which presents a similar hindering barrier to the methyl group rotation) is made in order to check the effect of the molecular environment of the methyl group on IVR phenomena.

aC. Lapouge and D. Cavagnat, J.Phys. Chem., to be published

bD. Cavagnat and L. Lespade,J.Chem.Phys.106,7946,(1997)